The spring oscillates to the other end. The amplitude is 3.5m, k = 1.2, and angular frequency is = 4.3. Solve the frequency, wavelength, wave speed, period, wave number, and direction.


1
Expert's answer
2021-11-30T18:16:15-0500

Wave number: "k=1.2"

angular frequency: "\omega=4.3"


Wavelength:


"\lambda=2\pi/k=5.2\text{ m}."

Frequency:


"f=\omega/2\pi=4.3/2\pi=0.68\text{ Hz}."



Wave speed = frequency x wavelength

"=0.68Hz \u00d7 5.2m = 3.536m\/s"


Period:"T=1/f=1.2\text{ s}."


Wave number ="0.68\/3.536=0.1923"


Direction :Moving right